# Continuum limit of bipartite lattices - The SSH model

### v1

@@ I Introduction
Recent advances have spurred interest in extending discrete models like SSH to continuous frameworks, facilitating exploration of topological phenomena in e.g. photonic crystals Naz2018 ; Wang2022 ; Tang2022 ; Gupta2022 , acoustic metamaterials Li2018 ; Zheng2019 ; Coutant2022 ; Liu2023 , and quantum field theories Dong2019 ; Salvo2024 . Among the same lines, several works have explored 1D continuous SSH analogues by modifying geometric or dynamical properties, focusing on phenomenological similarities or/and discrepancies Plekhanov2020 ; Fruchart2020 ; Mivehvar2021 ; Liu2021 ; Li2022 . Experimental platforms such as optical superlattices in cold-atom physics enable studies of topological properties in continuous systems Dalibard2018 ; Zhang2018 ; Cooper2019 ; Katsaris2024 .
@@
In the present work, we introduce a continuous 1D model that reproduces the infinite discrete SSH models’ bulk spectrum while supporting chiral zero-energy edge states when restricted to finite space. It is non-local, with a parameter specifying the spatial scale of non-locality in the model. This parameter allows for the study of the transition from the non-local to the local description, observing at the same time the associated modifications in the bulk spectrum and corresponding topological properties. In addition to being non-local, the model also differs from previously proposed Schrödinger-type models since it does not contain external potential terms. As a consequence, there are no additional energy bands, and the chiral symmetry is exactly preserved. Furthermore, the model differs from Dirac-type Hamiltonians because it is non-local, which turns out to be crucial for obtaining energy bands in the bulk, which are periodic functions of the wave vector $k$. When formulated on a finite space interval, the associated spectrum contains a flat band at zero energy involving an infinite countable set of exponentially localized edge states. This property suggests a reconsideration of bulk-boundary correspondence in the case of spatially continuous systems with non-trivial topology.
@@ II Non-local continuous SSH: spectrum and bulk topology
$\{\mathbf{H},\bm{\sigma}_{z}\}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [\mathbf{H},\bm{\sigma}_{x}\mathbf{\Pi}(x)]=0$ (2)
@@
$\mathbf{H}=\begin{pmatrix}0&w\leavevmode\nobreak\ \mathbf{T}(-a)+v\\ w\leavevmode\nobreak\ \mathbf{T}(a)+v&0\end{pmatrix}$ (4)
@@
$\gamma_{\pm}=\displaystyle{\int_{-\frac{\pi}{a}}^{\frac{\pi}{a}}}\mathrm{i}\langle u_{k}^{(\pm)}|\partial_{k}u_{k}^{(\pm)}\rangle\leavevmode\nobreak\ dk=\begin{cases}\phantom{\pm}0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{for}\leavevmode\nobreak\ \leavevmode\nobreak\ |v|>|w|\\ \pm\pi,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{for}\leavevmode\nobreak\ \leavevmode\nobreak\ |v|<|w|\end{cases}$ (14)
@@ III From local to non-local SSH in continuous space
In this section, our strategy is to expand the translation operator $\mathbf{T}(a)$ in Eq. (4) in powers of $a\frac{d}{dx}$ keeping terms up to order $O((a\frac{d}{dx})^{n})$. Thus, for increasing $n$, we obtain a more accurate local approximation $\mathbf{H}^{(n)}$ of the Hamiltonian $\mathbf{H}$ in Eq. (1). Clearly, in the limit $n\to\infty$ we recover the non-local Hamiltonian $\mathbf{H}$. In fact, as $n$ increases, higher and higher derivative terms appear in the off-diagonal elements of $\mathbf{H}^{(n)}$, making the corresponding approximate Hamiltonian less and less local. We will consider here the cases $n=0,\leavevmode\nobreak\ 1$ and $2$ focusing on the bulk spectrum and its topological characteristics. Notice that for each obtained model Hamiltonian $\mathbf{H}^{(n)}$ the associated spectrum is calculated exactly, without assuming that $a$ is a small parameter.
@@ IV Zero energy topological flat mini band in the finite non-local SSH model in continuous space
$v(x)=\begin{cases}v_{L}\leavevmode\nobreak\ \mathrm{for}\leavevmode\nobreak\ x\in[-{L\over 2},{L\over 2}]\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{for}\leavevmode\nobreak\ x\notin[-{L\over 2},{L\over 2}]\end{cases}$
$\mathrm{and}\leavevmode\nobreak\ \leavevmode\nobreak\ w(x)=\begin{cases}w_{L}\leavevmode\nobreak\ \mathrm{for}\leavevmode\nobreak\ x\in[-{L\over 2},{L\over 2}]\\ 0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \;\mathrm{for}\leavevmode\nobreak\ x\notin[-{L\over 2},{L\over 2}]\end{cases}$ (23)
Promoting the parameter $w$ to a function $w(x)$ introduces a subtle issue: the ordering of factors in $w\mathbf{T}(\pm a)$ in Eq. (4). It follows that, for the finite Hamiltonian $\mathbf{H}_{L}$ to be Hermitian, the translation operator must also act on the function $w(x)$ to its right, yielding $\mathbf{T}(\pm a)\left(w(x)..\right)$. This becomes evident upon considering the $O(a{d\over dx})$ expansion of $\mathbf{T}(\pm a)$ and imposing the Hermiticity of $\mathbf{H}^{(1)}$ in Eq. (15), under the assumption of a spatially varying $w(x)$. Taking this into account, the eigenvalue problem for the Hamiltonian $\mathbf{H}_{L}$ leads to the equations:
@@
$\psi_{s}(x)=U_{s}(x)e^{q_{s}x}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ s=A,\leavevmode\nobreak\ B\leavevmode\nobreak\ \leavevmode\nobreak\ ;U_{s}(x+a)=U_{s}(x)$ (25)
@@
$s=A,\leavevmode\nobreak\ B\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \eta_{A}=-1,\leavevmode\nobreak\ \eta_{B}=1$ (26)
@@
$U_{s}(x)=\cos\left(\frac{2\pi n_{s}}{a}x+\phi_{n_{s}}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ n_{s}\in\mathbb{N}$ (27)
where the phase $\phi_{n_{s}}$ is determined by the boundary conditions. For the assumed hard-wall boundary conditions we obtain:
$U_{s}\left(\displaystyle{-\frac{L}{2}}\right)=U_{s}\left(\displaystyle{\frac{L}{2}}\right)=0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ s=A,\leavevmode\nobreak\ B$ (28)
@@
$s=A,\leavevmode\nobreak\ B\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ n_{s}\in\mathbb{N},\leavevmode\nobreak\ m_{s}\in\mathbb{Z}$ (29)
As a consequence, Eq. (29) implies the existence of an infinite, countable set of eigenspinors to $\mathbf{H}_{L}$ classified by two integer numbers $(n_{s},m_{s})$ for each spinor component $\psi_{A}$, $\psi_{B}$. Both components are modulated by an exponential envelope leading to the localization to the left (right) in the interval ${L\over 2},{L\over 2}]$ of the component $\psi_{A}(x)$ ($\psi_{B}(x)$) respectively. In Fig. (4) we show the magnitudes $|\psi_{A}(x)|$, $|\psi_{B}(x)|$ for the zero energy eigenstate with $(n_{A},m_{A})=(3,1),\leavevmode\nobreak\ (n_{B},m_{B})=(5,2)$ on a logarithmic scale. A break on the $x$-axis is introduced to allow for better illustration of the functional behavior close to the boundaries.
Figure 4: The magnitude of the zero energy eigenstate-components $\psi_{s}(x)$ ($s=A$, $B$) with $(n_{A},m_{A})=(3,1),\leavevmode\nobreak\ (n_{B},m_{B})=(5,2)$ on a logarithmic scale. The parameter values used are $v_{0}=0.5$, $w_{0}=1$, $L=10$ and $a=0.2$ all in arbitrary units. The exponential envelope with characteristic exponent ${\frac{1}{a}\ln|\frac{w_{0}}{v_{0}}\Big{|}}\approx 1.5$ is clearly recognizable.
@@ V Concluding remarks
We introduce a continuous, infinitely extended, non-local model that captures nearly all the phenomenological features of the discrete SSH model. It shares the same bulk energy spectrum and bulk eigenstates. The associated Zak phase also behaves identically to that of the SSH model, indicating a similar topological structure. The model includes a parameter $a$, with dimension length, which represents a spatial scale and is responsible for the models’ non-locality. This parameter allows for a systematic study of the transition from a non-local to a local description. Notably, the energy spectrum and the corresponding eigenstates become exactly identical to those of the SSH model through appropriate tuning of this parameter.
@@
Our method for deriving continuous analogues of bipartite lattices is quite general and can be extended to other bipartite systems beyond SSH, as well as to more complex multipartite lattices. Finally, a note is in order concerning a corresponding experimental realization of our continuous non-local model. The standard SSH model has been implemented and analyzed experimentally in a variety of different physical situations and setups, such as the topological band structure of ultracold trapped atoms Cooper2019 , water waves in channels in the linear and nonlinear regime Anglart2025 , photonic superlattices Wang2021 ; Klauck2021 , and mechanical spring-connected systems Thatcher2022 . Non-locality could possibly be added to some of these setups analogous to the case of a helix, where long-range hopping across windings is enabled by the curved geometry Stockhofe2015 . Still, whether our spatially non-local continuous space analogues of bipartite (or multipartite) lattices can be realized experimentally remains an open and challenging question for future research.

### v2

@@ I Introduction
Recent advances have spurred interest in extending discrete models like SSH to continuous frameworks, facilitating exploration of topological phenomena in e.g. photonic crystals Naz2018 ; Wang2022 ; Tang2022 ; Gupta2022 , acoustic metamaterials Li2018 ; Zheng2019 ; Coutant2022 ; Liu2023 , and quantum field theories Dong2019 ; Salvo2024 . Among the same lines, several works have explored 1D continuous SSH analogues by modifying geometric or dynamical properties, focusing on phenomenological similarities or/and discrepancies Coutant2021 ; Lustig2017 ; Quemerais1993 ; Zoli04 ; Goren2018 . Experimental platforms such as optical superlattices in cold-atom physics enable studies of topological properties in continuous systems Dalibard2018 ; Zhang2018 ; Cooper2019 ; Katsaris2024 .
@@
In the present work, we investigate the possiblity to formulate an one-dimensional model defined in continuous space which combines the existence of finite Brillouin zone (infinite system) with the emergence of localized topologically protected edge modes (finite system). To this end, we introduce a continuous 1D model that reproduces the infinite discrete SSH models’ bulk spectrum while supporting chiral zero-energy edge states when restricted to finite space. It is non-local, with a parameter specifying the spatial scale of non-locality in the model. This parameter controls the period of the Brillouin zone as well as the localization length of the emerging edge states. Within this model it is possible to study the transition from non-local to local description, observing at the same time the associated modifications in the bulk spectrum and the corresponding topological properties. In addition to being non-local, the model also differs from previously proposed Schrödinger-type models since it does not contain external potential terms. As a consequence, there are no additional energy bands, and the chiral symmetry is exactly preserved. Furthermore, the model differs from Dirac-type Hamiltonians because it is non-local, which turns out to be crucial for obtaining energy bands in the bulk, which are periodic functions of the wave vector $k$. When formulated on a finite space interval, the associated spectrum contains a flat band at zero energy involving an infinite countable set of exponentially localized edge states. This property suggests a reconsideration of bulk-boundary correspondence in the case of spatially continuous systems with non-trivial topology.
@@ II Non-local continuous SSH: spectrum and bulk topology
$\{\mathbf{H},\bm{\sigma}_{z}\}=0\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ [\mathbf{H},\bm{\sigma}_{x}\mathbf{\Pi}(x)]=0$ (2)
@@
$\mathbf{H}=\begin{pmatrix}0&w\penalty 10000\ \mathbf{T}(-a)+v\\ w\penalty 10000\ \mathbf{T}(a)+v&0\end{pmatrix}$ (4)
@@
$\gamma_{\pm}=\displaystyle{\int_{-\frac{\pi}{a}}^{\frac{\pi}{a}}}\mathrm{i}\langle u_{k}^{(\pm)}|\partial_{k}u_{k}^{(\pm)}\rangle\penalty 10000\ dk=\begin{cases}\phantom{\pm}0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathrm{for}\penalty 10000\ \penalty 10000\ |v|>|w|\\ \pm\pi,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathrm{for}\penalty 10000\ \penalty 10000\ |v|<|w|\end{cases}$ (14)
@@ III From local to non-local SSH in continuous space
In this section, our strategy is to expand the translation operator $\mathbf{T}(a)$ in Eq. (4) in powers of $a\frac{d}{dx}$ keeping terms up to order $O((a\frac{d}{dx})^{n})$. Thus, for increasing $n$, we obtain a more accurate local approximation $\mathbf{H}^{(n)}$ of the Hamiltonian $\mathbf{H}$ in Eq. (1). Clearly, in the limit $n\to\infty$ we recover the non-local Hamiltonian $\mathbf{H}$. In fact, as $n$ increases, higher and higher derivative terms appear in the off-diagonal elements of $\mathbf{H}^{(n)}$, making the corresponding approximate Hamiltonian less and less local. We will consider here the cases $n=0,\penalty 10000\ 1$ and $2$ focusing on the bulk spectrum and its topological characteristics. Notice that for each obtained model Hamiltonian $\mathbf{H}^{(n)}$ the associated spectrum is calculated exactly, without assuming that $a$ is a small parameter.
@@ IV Zero energy topological flat mini band in the finite non-local SSH model in continuous space
$v(x)=\begin{cases}v_{L}\penalty 10000\ \mathrm{for}\penalty 10000\ x\in[-{L\over 2},{L\over 2}]\\ 0\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathrm{for}\penalty 10000\ x\notin[-{L\over 2},{L\over 2}]\end{cases}$
$\mathrm{and}\penalty 10000\ \penalty 10000\ w(x)=\begin{cases}w_{L}\penalty 10000\ \mathrm{for}\penalty 10000\ x\in[-{L\over 2},{L\over 2}]\\ 0\penalty 10000\ \penalty 10000\ \penalty 10000\ \;\mathrm{for}\penalty 10000\ x\notin[-{L\over 2},{L\over 2}]\end{cases}$ (23)
with $v_{L}$, $w_{L}$ constant. Promoting the parameter $w$ to a function $w(x)$ introduces a subtle issue: the ordering of factors in $w\mathbf{T}(\pm a)$ in Eq. (4). It follows that, for the finite Hamiltonian $\mathbf{H}_{L}$ to be Hermitian, the translation operator must also act on the function $w(x)$ to its right, yielding $\mathbf{T}(\pm a)\left(w(x)..\right)$. This becomes evident upon considering the $O(a{d\over dx})$ expansion of $\mathbf{T}(\pm a)$ and imposing the Hermiticity of $\mathbf{H}^{(1)}$ in Eq. (15), under the assumption of a spatially varying $w(x)$. Taking this into account, the eigenvalue problem for the Hamiltonian $\mathbf{H}_{L}$ leads to the equations:
@@
$\psi_{s}(x)=U_{s}(x)e^{q_{s}x}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ ,\penalty 10000\ s=A,\penalty 10000\ B\penalty 10000\ \penalty 10000\ ;U_{s}(x+a)=U_{s}(x)$ (25)
@@
$s=A,\penalty 10000\ B\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \eta_{A}=-1,\penalty 10000\ \eta_{B}=1$ (26)
@@
$U_{s}(x)=\cos\left(\frac{2\pi n_{s}}{a}x+\phi_{n_{s}}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ n_{s}\in\mathbb{N}$ (27)
where the phase $\phi_{n_{s}}$ is determined by the boundary conditions. It is also worth to notice here that the scale of non-locality $a$ influences both the value of the localization length characterizing the exponential decay as well as the period of the oscillating part of these zero modes.
For the assumed hard-wall boundary conditions we obtain:
$U_{s}\left(\displaystyle{-\frac{L}{2}}\right)=U_{s}\left(\displaystyle{\frac{L}{2}}\right)=0\penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ s=A,\penalty 10000\ B$ (28)
@@
$s=A,\penalty 10000\ B\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ n_{s}\in\mathbb{N},\penalty 10000\ m_{s}\in\mathbb{Z}$ (29)
As a consequence, Eq. (29) implies the existence of an infinite, countable set of eigenspinors to $\mathbf{H}_{L}$ classified by two integer numbers $(n_{s},m_{s})$ for each spinor component $\psi_{A}$, $\psi_{B}$. Both components are modulated by an exponential envelope leading to the localization to the left (right) in the interval ${L\over 2},{L\over 2}]$ of the component $\psi_{A}(x)$ ($\psi_{B}(x)$) respectively. In Fig. (4) we show the magnitudes $|\psi_{A}(x)|$, $|\psi_{B}(x)|$ for the zero energy eigenstate with $(n_{A},m_{A})=(3,1),\penalty 10000\ (n_{B},m_{B})=(5,2)$ on a logarithmic scale. A break on the $x$-axis is introduced to allow for better illustration of the functional behavior close to the boundaries.
Figure 4: The magnitude of the zero energy eigenstate-components $\psi_{s}(x)$ ($s=A$, $B$) with $(n_{A},m_{A})=(3,1),\penalty 10000\ (n_{B},m_{B})=(5,2)$ on a logarithmic scale. The parameter values used are $v_{0}=0.5$, $w_{0}=1$, $L=10$ and $a=0.2$ all in arbitrary units. The exponential envelope with characteristic exponent ${\frac{1}{a}\ln|\frac{w_{0}}{v_{0}}\Big|}\approx 1.5$ is clearly recognizable. Notice that $\psi_{s}(x)$ under the action of the operator $\bm{\sigma}_{x}\mathbf{\Pi}(x)$ (generalized parity) transforms to the function $\tilde{\psi}_{s}(x)$ with $\tilde{\psi}_{A}(x)=\psi_{B}(-x)$ and $\tilde{\psi}_{B}(x)=\psi_{A}(-x)$.
@@ V Concluding remarks
We introduce a continuous, infinitely extended, non-local model that captures nearly all the phenomenological features of the discrete SSH model. It shares the same bulk energy spectrum and bulk eigenstates. The associated Zak phase also behaves identically to that of the SSH model, indicating a similar topological structure. The model includes a parameter $a$, with dimension length, which represents a spatial scale and is responsible for the models’ non-locality. Notably, the energy spectrum and the corresponding eigenstates become exactly identical to those of the SSH model through appropriate tuning of this parameter.
@@
Our method for deriving continuous analogues of bipartite lattices is quite general and can be extended to other bipartite systems beyond SSH, as well as to more complex multipartite lattices. Furthermore, it would be interesting to consider the possibility to formulate the model in two spatial dimensions and explore the eventual emergence of localized states there.
Finally, a note is in order concerning a corresponding experimental realization of our continuous non-local model. The standard SSH model has been implemented and analyzed experimentally in a variety of different physical situations and setups, such as the topological band structure of ultracold trapped atoms Cooper2019 , water waves in channels in the linear and nonlinear regime Anglart2025 , photonic superlattices Wang2021 ; Klauck2021 , and mechanical spring-connected systems Thatcher2022 . Non-locality could possibly be added to some of these setups analogous to the case of a helix, where long-range hopping across windings is enabled by the curved geometry Stockhofe2015 . Still, whether our spatially non-local continuous space analogues of bipartite (or multipartite) lattices can be realized experimentally remains an open and challenging question for future research.